# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v1

@@ 3.4 Decomposing the resampling component
$V_{\mathrm{out}}(x):=\operatorname{Var}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta}\right]\right),$
which captures variability induced by the resampling design across trees. This component measures how the average prediction at $x$ changes across resampling realizations, before accounting for any additional variability from the remaining tree-level randomization. Conditioning on the resampling indicators and using the weighted representation of the tree predictor,
$\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta}\right]=\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta})\,Y_{i},\qquad\bar{W}_{i}(x;\mathcal{I}_{\theta}):=\mathbb{E}\!\left[W_{i}(x;\theta)\mid\mathcal{I}_{\theta}\right],$
where $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ is the expected weight assigned to observation $i$ at $x$ under the remaining tree-level randomization, given the resampling realization. With this notation,
$V_{\mathrm{out}}(x)=\operatorname{Var}\!\left(\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta})\,Y_{i}\right).$
Conceptually, $V_{\mathrm{out}}(x)$ isolates the effect of resampling on prediction. Different resampling realizations $\mathcal{I}_{\theta}$ change which observations are exposed to the tree, which in turn alters the distribution of the induced averaging set $A_{\theta}(x)$ and hence shifts the expected local averaging rule $\bar{W}(x;\mathcal{I}_{\theta})$. Unlike $V_{\mathrm{in}}(x)$, which reflects instability in how a given resampled dataset is locally partitioned, $V_{\mathrm{out}}(x)$ reflects variability arising from changes in which observations are even eligible to contribute to the prediction at $x$.
@@ 4.1 Law of Total Covariance Decomposition
We now examine the covariance term $C_{T}(x)$ appearing in Theorem 1. From the finite-sample variance identity, $C_{T}(x)$ is the limiting contribution to predictive variability as $B\to\infty$, representing irreducible dependence between tree-level prediction rules induced by the forest design at fixed sample size. Although the trees indexed by $\theta$ and $\theta^{\prime}$ are generated independently, their predictions at $x$ are generally dependent because both act on the same fixed dataset through randomized grouping mechanisms. To isolate the sources of this dependence, we condition on the resampling indicators $\mathcal{I}_{\theta}$ and $\mathcal{I}_{\theta^{\prime}}$, which determine which observations are exposed to each tree. Conditional on these indicators, any remaining dependence reflects how the forest construction induces related prediction rules at $x$ under independent tree generation.
Applying the law of total covariance with respect to the resampling indicators yields
$C_{T}(x)=\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}}\right)\right]+\operatorname{Cov}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta}\right],\mathbb{E}\!\left[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}}\right]\right).$ (3)
Because the two trees are generated independently, $\mathcal{I}_{\theta^{\prime}}$ carries no information about $T_{\theta}(x)$ beyond $\mathcal{I}_{\theta}$, and likewise with roles reversed; we therefore suppress redundant conditioning arguments implied by independence.
The two terms in (3) correspond to distinct design mechanisms. The first term captures dependence arising from joint inclusion of observations in the averaging sets $A_{\theta}(x)$ and $A_{\theta^{\prime}}(x)$, whereby the same outcomes receive positive weight in both trees. The second term captures dependence arising from alignment of the induced prediction rules at $x$: even when the averaging sets are disjoint, independently generated trees may induce similar local averaging behavior after conditioning on their respective resampling realizations.
This decomposition is exact at fixed sample size. Resampling schemes govern the first term through joint inclusion probabilities in the averaging sets, while feature-level randomization and split selection govern the second term through stability or alignment of the induced grouping mechanism around $x$. The following subsections analyze these components in turn.
@@ 4.2 Covariance Induced by Shared Training Observations
We analyze first the leading term $\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}}\right)\right]$, which captures dependence arising from reuse of the same training observations in the predictions at $x$. This mechanism is present whenever the same observation receives nonzero weight in both trees at $x$. Joint inclusion in the resampling realizations is necessary but not sufficient; the observation must also lie in the terminal region containing $x$ in both trees.
Using the weighted representation $T_{\theta}(x)=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}$ and conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$, the covariance expands as
$\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}}\right)$ $=\operatorname{Cov}\!\left(\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i},\;\sum_{j=1}^{n}W_{j}(x;\theta^{\prime})\,Y_{j}\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}}\right)$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}(x;\theta)\,W_{j}(x;\theta^{\prime})\,\operatorname{Cov}(Y_{i},Y_{j}\mid X_{i},X_{j}).$
Under conditional independence of outcomes across observational units given the fixed design points, $\operatorname{Cov}(Y_{i},Y_{j}\mid X_{i},X_{j})=0$ for $i\neq j$, so the double sum reduces to
$\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}}\right)=\sum_{i=1}^{n}W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\,\sigma_{i}^{2}.$
Where $\sigma_{i}^{2}=\operatorname{Var}(Y|X=X_{i})$ as defined in Section 3.3. Taking expectations over the tree-generating randomization yields
$\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}}\right)\right]=\sum_{i=1}^{n}\mathbb{E}\!\left[W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\right]\sigma_{i}^{2}.$
@@
$\mathbb{E}\!\left[W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\right]=\mathbb{P}\!\left(I_{\theta,i}=1,\;I_{\theta^{\prime},i}=1\right)\,\mathbb{E}\!\left[W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\mid I_{\theta,i}=I_{\theta^{\prime},i}=1\right],$
@@ 4.3 Covariance Induced by Partition Alignment
$\operatorname{Cov}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta}\right],\mathbb{E}\!\left[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}}\right]\right),$
which captures dependence arising from alignment of the induced local geometry around $x$ across independently generated trees. This mechanism does not require any observation to receive positive weight in both trees at $x$ in a given realization. Instead, it arises when distinct trees, potentially trained on disjoint exposed subsets, induce the same (or very similar) terminal region containing $x$ and therefore apply similar local averaging rules. Concretely, two trees trained on different observations may split on the same variables in the approximately the same order at similar thresholds in a neighborhood of $x$; in that case $x$ is routed into essentially the same geometric terminal node, and each tree averages outcomes from the same covariate-defined subpopulation even if the contributing training observations are disjoint.
@@
Conditional on the resampling indicators $\mathcal{I}_{\theta}$, the remaining randomness in $T_{\theta}(x)$ arises from feature subsampling and split selection, which determine the realized region $R_{\theta}(x)$ and hence the induced averaging set. Averaging over this remaining randomness induces the resampling-conditional prediction rule
$Z_{\theta}(x):=\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta}],\qquad Z_{\theta^{\prime}}(x)=\mathbb{E}[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}}].$
@@
$\bar{W}_{i}(x;\mathcal{I}_{\theta}):=\mathbb{E}\!\left[W_{i}(x;\theta)\;\middle|\;\mathcal{I}_{\theta}\right],\qquad i=1,\ldots,n,$
@@
The mean weights $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ do not correspond to membership in a single realized terminal region. Rather, they average over the distribution of terminal regions induced by split-level randomization, assigning observation $i$ mass proportional to how frequently it is grouped with $x$ across possible tree realizations. This resampling-conditional mean weight is an intermediate quantity between the realized tree-level weight $W_{i}(x;\theta)$ and the forest-level expected weight $\mathbb{E}[W_{i}(x;\theta)]$; its introduction is what makes the alignment mechanism formally identifiable as a distinct component of the covariance structure. In this sense, $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ represents the resampling-conditional local neighborhood around $x$ induced by the procedure.
@@
$C_{\mathrm{align}}(x):=\operatorname{Cov}\!\left(Z_{\theta}(x),Z_{\theta^{\prime}}(x)\right),$
where the covariance is taken under the design-based distribution induced by the independent resampling indicators $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$. Expanding yields
$C_{\mathrm{align}}(x)=\sum_{i=1}^{n}\sum_{j=1}^{n}\operatorname{Cov}(Y_{i},Y_{j}\mid X_{i},X_{j})\,\mathbb{E}\!\left[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{j}(x;\mathcal{I}_{\theta^{\prime}})\right].$
Under conditional independence of outcomes across observational units given the fixed design points, $\operatorname{Cov}(Y_{i},Y_{j}\mid X_{i},X_{j})=0$ for $i\neq j$, and therefore
$C_{\mathrm{align}}(x)=\sum_{i=1}^{n}\sigma_{i}^{2}\;\mathbb{E}\!\left[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})\right].$ (4)
Expression (4) makes the alignment mechanism explicit in design-based terms. The joint expectation $\mathbb{E}[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})]$ measures whether independent trees tend, on average, to place predictive mass near $x$ on the same local subset of design points after averaging over split-level randomization. When routing near $x$ is stable across tree realizations—so that similar split conditions recur and $x$ repeatedly falls into essentially the same terminal region—the procedure concentrates on a restricted family of local regions in covariate space. The observations populating those regions, though potentially disjoint across trees, are drawn from the same conditional subpopulation defined by the shared partition geometry. Both tree predictions therefore estimate the same conditional mean, yielding positive covariance even in the absence of shared training observations.

### v2

@@ 3.4 Decomposing the resampling component
$V_{\mathrm{out}}(x):=\operatorname{Var}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right]\;\middle|\;X\right),$
which captures variability induced by the resampling design across trees, conditional on the realized covariate configuration $X$. This component measures how the average prediction at $x$ changes across resampling realizations, before accounting for any additional variability from the remaining tree-level randomization.
Conditioning on the resampling indicators and using the weighted representation of the tree predictor,
$\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right]=\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta})\,Y_{i},\qquad\bar{W}_{i}(x;\mathcal{I}_{\theta}):=\mathbb{E}\!\left[W_{i}(x;\theta)\mid\mathcal{I}_{\theta},X\right],$
where $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ is the expected weight assigned to observation $i$ at $x$ under the remaining tree-level randomization, given the resampling realization and the fixed covariate configuration. With this notation,
$V_{\mathrm{out}}(x)=\operatorname{Var}\!\left(\sum_{i=1}^{n}\bar{W}_{i}(x;\mathcal{I}_{\theta})\,Y_{i}\;\middle|\;X\right).$
Conceptually, $V_{\mathrm{out}}(x)$ isolates the effect of resampling on prediction under the design-based distribution conditional on $X$. Different resampling realizations $\mathcal{I}_{\theta}$ change which observations are exposed to the tree, which in turn alters the distribution of the induced averaging set $A_{\theta}(x)$ and hence shifts the expected local averaging rule $\bar{W}(x;\mathcal{I}_{\theta})$. Unlike $V_{\mathrm{in}}(x)$, which reflects instability in how a given resampled dataset is locally partitioned, $V_{\mathrm{out}}(x)$ reflects variability arising from changes in which observations are even eligible to contribute to the prediction at $x$.
@@ 4.1 Law of Total Covariance Decomposition
We now examine the covariance term $C_{T}(x)$ appearing in Theorem 1. From the finite-sample variance identity, $C_{T}(x)$ is the limiting contribution to predictive variability as $B\to\infty$, representing irreducible dependence between tree-level prediction rules induced by the forest design at fixed sample size. Although the trees indexed by $\theta$ and $\theta^{\prime}$ are generated independently, their predictions at $x$ are generally dependent because both act on the same realized outcomes under the fixed covariate configuration $X$ through randomized grouping mechanisms. To isolate the sources of this dependence, we condition on the resampling indicators $\mathcal{I}_{\theta}$ and $\mathcal{I}_{\theta^{\prime}}$, which determine which observations are exposed to each tree. Conditional on these indicators and on $X$, any remaining dependence reflects how the forest construction induces related prediction rules at $x$ under independent tree generation.
Applying the law of total covariance conditional on $X$ with respect to the resampling indicators yields
$C_{T}(x)=\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)\;\middle|\;X\right]+\operatorname{Cov}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right],\mathbb{E}\!\left[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}},X\right]\;\middle|\;X\right).$ (3)
Because the two trees are generated independently, $\mathcal{I}_{\theta^{\prime}}$ carries no information about $T_{\theta}(x)$ beyond $\mathcal{I}_{\theta}$ once $X$ is fixed, and likewise with roles reversed; we therefore suppress redundant conditioning arguments implied by independence.
The two terms in (3) correspond to distinct design mechanisms. The first term captures dependence arising from joint inclusion of observations in the averaging sets $A_{\theta}(x)$ and $A_{\theta^{\prime}}(x)$, whereby the same outcomes receive positive weight in both trees. The second term captures dependence arising from alignment of the induced prediction rules at $x$: even when the averaging sets are disjoint, independently generated trees may induce similar local averaging behavior after conditioning on their respective resampling realizations and the fixed covariate configuration.
This decomposition is exact at fixed sample size and fixed $X$. All expectations and covariances are taken with respect to the joint randomness of $(Y\mid X)$ and the tree-generating mechanism $\theta$. Resampling schemes govern the first term through joint inclusion probabilities in the averaging sets, while feature-level randomization and split selection govern the second term through stability or alignment of the induced grouping mechanism around $x$. The following subsections analyze these components in turn.
@@ 4.2 Covariance Induced by Shared Training Observations
We first analyze the leading term, $\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)\;\middle|\;X\right]$, which captures dependence arising from reuse of the same training observations in the predictions at $x$. This mechanism is present whenever the same observation receives nonzero weight in both trees at $x$. Joint inclusion in the resampling realizations is necessary but not sufficient; the observation must also lie in the terminal region containing $x$ in both trees.
Using the weighted representation $T_{\theta}(x)=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}$ and conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X)$, the covariance expands as
$\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)$ $=\operatorname{Cov}\!\left(\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i},\;\sum_{j=1}^{n}W_{j}(x;\theta^{\prime})\,Y_{j}\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}(x;\theta)\,W_{j}(x;\theta^{\prime})\,\operatorname{Cov}(Y_{i},Y_{j}\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X).$
Because the resampling indicators are drawn independently of the outcome-generating process, conditioning on $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$ does not alter the joint distribution of $(Y_{i},Y_{j})$ given $X$. Moreover, under conditional independence of outcomes across observational units given the fixed covariate configuration, $\operatorname{Cov}(Y_{i},Y_{j}\mid X)=0$ for $i\neq j$, so the double sum reduces to
$\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\;\middle|\;\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)=\sum_{i=1}^{n}W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\,\sigma_{i}^{2},$
where we define
$\sigma_{i}^{2}:=\operatorname{Var}(Y_{i}\mid X),$
the conditional outcome variance at the realized design point $X_{i}$. This quantity is a property of the fixed-design outcome model and does not depend on the forest construction. Taking expectations over the tree-generating randomization conditional on $X$ yields
$\mathbb{E}\!\left[\operatorname{Cov}\!\left(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}},X\right)\;\middle|\;X\right]=\sum_{i=1}^{n}\mathbb{E}\!\left[W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\;\middle|\;X\right]\sigma_{i}^{2}.$
@@
$\mathbb{E}\!\left[W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\;\middle|\;X\right]=\mathbb{P}\!\left(I_{\theta,i}=1,\;I_{\theta^{\prime},i}=1\right)\,\mathbb{E}\!\left[W_{i}(x;\theta)\,W_{i}(x;\theta^{\prime})\mid I_{\theta,i}=I_{\theta^{\prime},i}=1,X\right],$
@@ 4.3 Covariance Induced by Partition Alignment
$\operatorname{Cov}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right],\mathbb{E}\!\left[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}},X\right]\;\middle|\;X\right),$
which captures dependence arising from alignment of the induced local geometry around $x$ across independently generated trees. This mechanism does not require any observation to receive positive weight in both trees at $x$ in a given realization. Instead, it arises when distinct trees, potentially trained on disjoint exposed subsets, induce the same (or very similar) terminal region containing $x$ and therefore apply similar local averaging rules. Concretely, two trees trained on different observations may split on the same variables in approximately the same order at similar thresholds in a neighborhood of $x$; in that case $x$ is routed into essentially the same geometric terminal node, and each tree averages outcomes from the same covariate-defined subpopulation even if the contributing training observations are disjoint.
@@
Conditional on the resampling indicators and fixed $X$, the remaining randomness in $T_{\theta}(x)$ arises from feature subsampling and split selection, which determine the realized region $R_{\theta}(x)$ and hence the induced averaging set. Averaging over this remaining randomness induces the resampling-conditional prediction rule
$Z_{\theta}(x):=\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},X],\qquad Z_{\theta^{\prime}}(x)=\mathbb{E}[T_{\theta^{\prime}}(x)\mid\mathcal{I}_{\theta^{\prime}},X].$
@@
$\bar{W}_{i}(x;\mathcal{I}_{\theta}):=\mathbb{E}\!\left[W_{i}(x;\theta)\;\middle|\;\mathcal{I}_{\theta},X\right],\qquad i=1,\ldots,n,$
@@
The mean weights $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ do not correspond to membership in a single realized terminal region. Rather, they average over the distribution of terminal regions induced by split-level randomization, assigning observation $i$ mass proportional to how frequently it is grouped with $x$ across possible tree realizations. This resampling-conditional mean weight is an intermediate quantity between the realized tree-level weight $W_{i}(x;\theta)$ and the forest-level expected weight $\mathbb{E}[W_{i}(x;\theta)\mid X]$; its introduction makes the alignment mechanism formally identifiable as a distinct component of the covariance structure. In this sense, $\bar{W}_{i}(x;\mathcal{I}_{\theta})$ represents the resampling-conditional local neighborhood around $x$ induced by the procedure.
@@
$C_{\mathrm{align}}(x):=\operatorname{Cov}\!\left(Z_{\theta}(x),Z_{\theta^{\prime}}(x)\;\middle|\;X\right),$
where the covariance is taken under the joint randomness of $(Y\mid X)$ and the independent resampling indicators $(\mathcal{I}_{\theta},\mathcal{I}_{\theta^{\prime}})$. Expanding yields
$C_{\mathrm{align}}(x)=\sum_{i=1}^{n}\sum_{j=1}^{n}\operatorname{Cov}(Y_{i},Y_{j}\mid X)\,\mathbb{E}\!\left[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{j}(x;\mathcal{I}_{\theta^{\prime}})\;\middle|\;X\right].$
Under conditional independence of outcomes across observational units given $X$, $\operatorname{Cov}(Y_{i},Y_{j}\mid X)=0$ for $i\neq j$, and therefore
$C_{\mathrm{align}}(x)=\sum_{i=1}^{n}\sigma_{i}^{2}\;\mathbb{E}\!\left[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})\;\middle|\;X\right],$ (4)
where $\sigma_{i}^{2}=\operatorname{Var}(Y_{i}\mid X)$.
Expression (4) makes the alignment mechanism explicit in design-based terms. The joint expectation $\mathbb{E}[\bar{W}_{i}(x;\mathcal{I}_{\theta})\,\bar{W}_{i}(x;\mathcal{I}_{\theta^{\prime}})\mid X]$ measures whether independent trees tend, on average, to place predictive mass near $x$ on the same local subset of design points after averaging over split-level randomization. When routing near $x$ is stable across tree realizations—so that similar split conditions recur and $x$ repeatedly falls into essentially the same terminal region—the procedure concentrates on a restricted family of local regions in covariate space. The observations populating those regions, though potentially disjoint across trees, are drawn from the same conditional subpopulation defined by the shared partition geometry. Both tree predictions therefore estimate the same conditional mean, yielding positive covariance even in the absence of shared training observations.